OpenStudy (anonymous):

write an equation for each line that contains the given pair of points (-6,1) (2,3) PLEASE SHOW ALL STEPS AND EXPLAIN

OpenStudy (anonymous):

1. Plug both coordinates into the slope formula: \[y _{2}-y _{1}/x _{2}-x _{1}\] to get 2/8, or 1/4. m = 1/4 = slope. 2. Plug your slope and either coordinate point (I'll use (2,3)) into the point-slope formula:\[y-y _{1}=m(x-x_{1})\] to get \[y-3=1/4(x-2)\]\[y-3=(1/4)x-(1/2)\]\[y=(1/4)x-(7/2)\]

OpenStudy (anonymous):

Yeah I do NOT understand that!!!! :/

OpenStudy (anonymous):

Plug both of your coordinates into the slope formula which I gave you. That will give you your slope (denoted by the variable 'm'). Then plug your newly found slope, along with either one of the coordinates (its your choice which one, but its obviously easier to use smaller coordinate), into the point-slope formula (plug in the 'm' you just found, as well as your coordinate (in this case, I chose (2,3) to be (x1,y1)). Then, just solve for y.

OpenStudy (anonymous):

my teacher found the slope for me like this \[m=\Delta y/\Delta x \] \[3-1/2-(-6)= 2/8=1/4\] and then I plugged in (2,3) and did this \[3=1/4(2) + b\] \[3(3) =1/2(3) =b\] \[9=3/2=b\] Now I'm stumped

OpenStudy (anonymous):

Ah, our slope formulas are the same. So, the equation for a line is:\[y=mx+b\] So, now that you found 'm' and 'b', plug it back into the slope-intercept formula. So, your new line would be:\[y=1/4x+3/2\]